# Rectangular recurrence relations in $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ invariant integrable models

### v1

@@ Acknowledgement
The proof of this theorem follows the method described in the appendix A. The starting point is the simple recurrence relation (4.6) for the Bethe vector ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-1})$. Applying the zero mode operator ${\sf T}_{n,n-1}$ to this relation and using the commutation relations between the zero modes and the monodromy entries (4.7) as well as the action of the zero mode operators on the off-shell Bethe vectors (4.8) we get a relation which involves terms proportional to $\chi_{n}$ and terms proportional to $\chi_{n-1}$. Since the relation involves only Bethe vectors, monodromy matrix entries $T_{ij}(z)$ and eigenvalues $\lambda_{i}(z)$, due to remark 2.2, the coefficients of these two independent twisting parameters yield two recurrence relations, one for the Bethe vectors ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{u}^{n-1})=\mathcal{Z}^{n}_{-n+1}\cdot{\mathbb{B}}({\bar{t}})$ and one for the Bethe vector ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{v}^{n-1})=\mathcal{Z}^{n-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})$. These relations are
@@
Referring to the calculations presented in appendix A for more details, we describe below the different ranges of the indices $\ell$, $k$ and the corresponding identities which should be used at each of the inductive step. Let us divide the whole domain $-n\leq\ell<k\leq n$ of the values of the indices $\ell$ and $k$ into three subdomains: $1\leq-\ell,k\leq n$, $0\leq\ell<k\leq n$ and $-n\leq\ell<k\leq 0$.

### v2

@@ Acknowledgement
The proof of this theorem follows the method described in the appendix A. The starting point is the simple recurrence relation (4.6) for the Bethe vector ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-1})$.
The recurrence relations (4.21) become the relations (4.31), (4.32), and (4.33) for $n=1$. They were proved in [21]. To prove the theorem 4.1 it is sufficient to consider the cases when $n>1$.
Applying the zero mode operator ${\sf T}_{n,n-1}$ to (4.6) and using the commutation relations between the zero modes and the monodromy entries (4.7) as well as the action of the zero mode operators on the off-shell Bethe vectors (4.8) we get a relation which involves terms proportional to $\chi_{n}$ and terms proportional to $\chi_{n-1}$. Since the relation involves only Bethe vectors, monodromy matrix entries $T_{ij}(z)$ and eigenvalues $\lambda_{i}(z)$, due to remark 2.2, the coefficients of these two independent twisting parameters yield two recurrence relations, one for the Bethe vectors ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{u}^{n-1})=\mathcal{Z}^{n}_{-n+1}\cdot{\mathbb{B}}({\bar{t}})$ and one for the Bethe vector ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{v}^{n-1})=\mathcal{Z}^{n-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})$. These relations are
@@
[…]